OpenStudy (anonymous):

If a ball is thrown vertically upward from the roof of 48 foot building with a velocity of 48 ft/sec, its height after seconds is s(t)=48+48t-16t^2. I got: What is the maximum height the ball reaches? 84feet But I need help on... What is the velocity of the ball when it hits the ground (height 0)?

OpenStudy (anonymous):

\[s(t) = 48+48t - 16t^2\\ s(t) = 48+ 32t \\32t = 48\]

OpenStudy (anonymous):

3/2?

OpenStudy (anonymous):

uhm, no 48/32 is 3/2?

OpenStudy (anonymous):